From GlueXWiki
Jump to: navigation, search

June 27, 2013 FDC meeting


  1. Production
  2. Engineering [9], [10] (Bill)
  3. Electronics (Chris, Nick)
  4. Full electronics test FDC E-log (Beni)
  5. Other


Participants: Bill, Dave, Chris, Nick, Simon, Vlad, Beni, Eugene, and Lubomir.


- Bill: two spacers are ready, the last one will be glued right after the meeting. Pictures of the spacers, the fixture and the gluing are attached above. The carbon fiber rods are glued to the g10 cylinder. The two gusset rings needed for this procedure were taken from the spare package which is sitting now on a granite table with two Lexan sheets on both sides. Tomorrow after the last spacer is ready, we will put the gussets back. This operation requires at least 4 people, since we use metal pins inserted in the six places when lifting the Lexan sheets. As in happened already when taking the gusset ring in the same way, the pin can lift some of the planes/cells up.

- Lubomir and Vlad showed some results from the thick cell study with a scope. We used two differential probes connected to wires from the two top cells. We used a scope as a trigger and fADC: we triggered on coincidences from the two wires, thus selecting almost vertical MIP that crossed the whole volume above the wires. We recorded manually about a thousand events and the plots linked above represent the results from their analysis. We increased the HV to +2450/-500V which corresponds to about 2-3*10^5 and it was stable for the whole weekend. With such high gain and very low noise (threshold of 5mV) we were able to see individual electrons with about 80% efficiency. There was 30-40 % loss of electrons (somehow different for the two cells) due to recombination (oxygen was 90-100 ppm) over the very long drift time of >20usec. Increasing the flow will reduce this effect. One can see the correlation in the cluster number counted by the two cells. The width of the difference agrees with the statistics. Interesting studies of the amplitudes of the single electrons show a nice exponential distribution of the gas gain with a slope that agrees to the average gas gain.


- Bill explained how to install the trigger scintillators on the FDC installation cart as shown in the 3D models above.

- We discussed the plan for the coming two months: finish with the spacers (end of this week) and the installation of the manifolds (next week) ; wait for the fiducial holders (end of next week) and then start installation of packages #1 and #2 on the cart with a spacer in between. After the third package is back from testing continue with the installation of packages #3 and #4; install the fiducials; make the gas connections between the packages (Bill still needs to find the right manifolds for the gas). At about middle of July, after all four packages are connected move them outside of the clean room and start practicing with the mock-up, connecting with cables one quadrant. Then, beginning of August move the packages to 126 and do the tests with the HV.

- At the same time we should start moving the equipment out of Blue Crab. According to Dave Tim is working to find some space for us in the TDEF building. Eugene: Rolf Ent stated that physics wants to keep the clean room. There are rumors that Hall C wants to disassemble (Walter is coordinating this) it and store it somewhere for later use, but nothing official. In any case we would like to have the right to use some area of the clean room if we need to do something on the FDC.


- Chris and Nick will collect all the signal cables for each package together: some of the cables are with Tom, that's why the cables at Blue Crab to be used for the mock-up are mixed.

- We may need contact cleaner for fixing the cathode pre-amp contacts; Chris will look for such cleaner.

Full electronics tests

- Now the whole package is equipped with electronics and all the cables connected.

- Beni showed results of the 4 cell tests. In the E-log entry 112 most of the pictures show the wire position as reconstructed by the strips. Interestingly, the width of the wire varies with the position in the chamber. It is small in the middle, then getting wider, then again small at the periphery. People suggested different explanations, but all this requires more studies, like plots of the widths along one strip, increasing the high voltage to see if the effect remains with higher signal and better strip resolution. Such variations resemble the variations we had of the charge from the strips due to frame deformation, but this is unlikely (Bill) since we have a strong-back now. Other than that, the noise on the strips and wires is as low as it was in the previous test of the packages partially equipped with electronics. Several cathode channel still require "fixing" by re-seating the pre-amps and eventually bending the cooling tubes.

- Plans: after fixing all the channels we would like to take data also at different high voltages; expect to be ready with the tests by July 8th in which week we want to move the package to Blue Crab. There will be no time now to test the new F1TDCs, since this will require new crate and with a ROC and TI, and also development of the firmware and software. Rather than that, we decided to test the new F1TDCs with the spare package that will be moved to 126 at the same week of July 8th.


- No meeting next week: 4th of July.